OpenStudy (anonymous):

Which of the following is a foci of the conic section with the equation x^2/16-y^2/4=1 (2squareroot5,0) (0,2squareroot5) (2squareroot3,0)

OpenStudy (mertsj):

I see this is exactly like the previous problem you posted. Would you like to learn to solve these problems yourself?

OpenStudy (anonymous):

yes !

OpenStudy (mertsj):

The foci are c units from the center do you know that?

OpenStudy (anonymous):

no..?

OpenStudy (mertsj):

Well, you know that now. It is always the case. Then you have to know whether the hyperbola opens up and down or right and left. Which does this do?

OpenStudy (mertsj):

Ok. let's approach it a slightly different way: \[\frac{(x-h)^2}{a^2}-\frac{(y-k)^2}{b^2}=1\]

OpenStudy (mertsj):

This is the equation of a hyperbola that opens to the right and left because the x^2 term is positive.

OpenStudy (mertsj):

The center is (h,k)

OpenStudy (mertsj):

a^2+b^2=c^2

OpenStudy (mertsj):

Now can you find your foci?

OpenStudy (anonymous):

okay i know that !

OpenStudy (mertsj):

So you can find c?

OpenStudy (mertsj):

And the center?

OpenStudy (anonymous):

so it would be A !

OpenStudy (anonymous):

right or wrong ?

OpenStudy (mertsj):

Very good!!